I have no clue what half-again is, the teacher never discussed it, please help. Here is the problem: The longest side of a triangle is half-again as long as the shortest side. The other side is 2 feet shorter than the longest side. The perimeter of the triangle is 22 feet. Find the length of each side of the triangle.
Answer by stanbon(48558)

The longest side of a triangle is half-again as long as the shortest side. The other side is 2 feet shorter than the longest side. The perimeter of the triangle is 22 feet. Find the length of each side of the triangle.
----------
Comment: half-again as long as 6 would be 6+(1/2)6 = 6+3 = 9
-------------------
Let the shortest side have length = x.
The longest side will be x+(1/2)x = (3/2)x
The "other side" is (3/2)x-2
-------------------
Perimeter = side + side + side
22 = x + (3/2)x + (3/2)x-2
24 = 4x
x = 6 feet (shortest side)
longest side = (3/2)x = 9 ft.
other side = 7 ft
================
Cheers,
Stan H.